y′′−8y′+16y=24x+2


1
Expert's answer
2021-06-16T06:20:41-0400

Homogeneous differential equation,

y′′−8y′+16y=0

Auxiliary equation,

m2-8m+16=0

(m-4)2=0

m=4,4 (repeated roots)

Therefore, solution is y=(a+bx)e4x.

Now, particular solution is given by,

"y_p=\frac{1}{(D-4)^2}24x+2\\
=\frac{1}{16}[1-\frac{D}{4}]^{-2}(24x+2)\\
=\frac{1}{16}[1+\frac{D}{2}+3\frac{D^2}{16}+•••
](24x+2)\\
=\frac{1}{16}[24x+2+12]\\
=\frac{1}{16}[24x+14]\\
=\frac{12x+7}{8}"

Thus, solution is "y=(a+bx)e^{4x}+\frac{12x+7}{8}"